### v2

# Discovery of Seven Cold and Distant Brown Dwarfs with JWST RUBIES

Abstract
We report near-infrared spectral model fits to seven distant L- and T-type dwarfs observed with the JWST Near Infrared Spectrograph (NIRSpec) as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES). Comparison of 0.9–2.5 m near-infrared spectra of these sources to spectral standards indicates spectral types spanning L1 to T8 and spectrophotometric distances spanning 800–3,000 pc. Fits to three grids of spectral models yield atmosphere parameters and spectrophotometric distances largely consistent with our classifications, although fits to L dwarf spectra indicate missing components to the models. Three of our sources have vertical displacements from the Galactic plane exceeding 1 kpc, and have high probabilities of membership in the Galactic thick disk population. Of these, the L dwarf RUBIES-BD-3 (RUBIES-EGS-3081) is well-matched to subdwarf standards, while the early T dwarf RUBIES-BD-5 (RUBIES-UDS-170428) is best fit by metal-poor atmosphere models; both be a thick disk or halo brown dwarfs. We critically examine the 1–5 m spectra of the current sample of 1–2 kpc mid- and late-T dwarfs, finding that temperature, surface gravity, metallicity, and vertical mixing efficiency can all contribute to observed variations in near-infrared spectral structure and the strength of the 4.2 m CO2 band. This work aims to guide ongoing JWST, Euclid, and other space-based spectral surveys that are expected to uncover thousands of low-temperature stars and brown dwarfs throughout the Milky Way.
show]aburgasser@ucsd.edu
1 Introduction
Brown dwarfs are the lowest-temperature and lowest-mass “stars” in the Milky Way, with insufficient mass (0.075 M⊙) to sustain core hydrogen fusion (1962AJ.....67S.579K; 1963ApJ...137.1121K; 1963PThPh..30..460H). While comprising a significant fraction of star-like objects in the Galactic disk (20%; Reyl__2018; Kirkpatrick_2021), these objects cool and dim over time, becoming exceedingly faint ( 18, 10; 2017AJ....154..147D). As such, the majority of low-temperature stars and brown dwarfs—classified as late-M, L, T and Y dwarfs—are detected in the immediate solar neighborhood (100 pc), and have predominantly solar elemental compositions. Only a small fraction of local dwarfs are metal-poor members of the Galactic thick disk (12%) and halo populations (0.5%; 2008ApJ...673..864J; 2020ApJ...898...77S). While rare, these long-lived sources are essential for understanding how metallicity influences the formation, evolution, and hydrogen-burning mass threshold brown dwarfs, and can potentially trace dynamical evolution and chemical enrichment throughout the history of the Milky Way (2004ApJS..155..191B; 2025ApJ...985...48H).
The high infrared sensitivity of the current generation of space-based telescopes, including JWST (2023PASP..135d8001R) and Euclid (2011arXiv1110.3193L), has made it possible to identify and study brown dwarfs at considerably larger distances. Building on tools developed for deep imaging and spectroscopic surveys conducted with the Hubble Space Telescope (2009ApJ...695.1591P; 2011ApJ...739...83R; 2016MNRAS.458..425V; 2022ApJ...924..114A), deep JWST surveys are now uncovering individual low-temperature stars and brown dwarfs beyond kiloparsec distances (2023ApJ...942L..29N; 2023ApJ...957L..27L; 2023MNRAS.523.4534W; 2024ApJ...962..177B; 2024MNRAS.529.1067H; 2024ApJ...964...66H; 2024ApJ...975...31H; 2025ApJ...980..230T; 2025arXiv251002026T; 2025arXiv251000111H). Early Euclid studies are uncovering hundreds of cool dwarfs to similar scales (2025arXiv250322497V; 2025arXiv250322559M; 2025ApJ...991...84D). Those surveys conducted at high galactic latitudes contain considerably higher fractions of thick disk and halo sources than in the local population, enabling more efficient and comprehensive studies of metal-poor low-mass stars and brown dwarfs.
In this article, we present the identification and spectral analysis of seven L and T dwarfs identified in deep spectroscopic data collected as part of the Red Unknowns: Bright Infrared Extragalactic Survey (RUBIES; 2025A&A...697A.189D). In Section 2, we describe the design and data associated with the RUBIES survey, and our method for identifying brown dwarfs from the spectral data. In Section 3, we present an empirical analysis of the spectra, including template classification and spectrophotometric distance estimates. In Section 4, we fit the spectra to three sets of model grids to obtain robust estimates of atmospheric parameters and their uncertainties, and evaluate the accuracy of the different grids over the broad range of specral types in our sample. In Section 5, we quantify the probabilities of membership in the Galactic thin disk, thick disk, and halo populations, and critically assess the source of spectral variations among mid- and late-type T dwarfs in our sample. We summarize our results in Section 6.
2 Sample
2.1 RUBIES Survey Data
RUBIES is a spectroscopic survey conducted with the JWST Near Infrared Spectrograph (NIRSpec; Birkmann_2022). The survey is focused on detecting and characterizing massive galaxies at high redshifts, drawing on sources selected from JWST Near Infrared Camera (NIRCam) imaging data obtained in the Extended Groth Strip (EGS) by the Cosmic Evolution Early Release Science Survey (CEERS, JWST-ERS-1345, PI S. Finkelstein), and in the Ultra Deep Survey (UDS) by the Public Release IMaging for Extragalactic Research Survey (PRIMER, JWST-GO-1837, PI J. Dunlap). RUBIES prioritizes very red sources, and thus has a higher chance of observing brown dwarfs than surveys with differing color selection. All RUBIES sources were acquired using the NIRSpec micro-shutter array (MSA; 2022A&A...661A..81F) with a three-shutter slitlet, with total integrations of 48 minutes acquired in each of the low-resolution Prism/Clear mode (0.6–5.3 m, = 30–300) and the medium-resolution G395M/F290LP mode (2.9–5.3 m, = 1000), using a three-point nodding pattern. Spectra were reduced using the JWST calibration pipeline version 1.12.5 (bushouse_2023_10022973) and msaexp (brammer_2023_8319596) as described in detail in 2025A&A...693A..60H and 2025A&A...697A.189D. Here, we focus on the NIRSpec/Prism data which provide the highest signal-to-noise (S/N) for our faint targets.
2.2 Selection of Brown Dwarfs
| Name | BD-1 | BD-2 | BD-3 | BD-4 | BD-5 | BD-6 | BD-7 |
| Survey Information | |||||||
| RUBIES ID | RUBIES-EGS-944743 | RUBIES-EGS-35616 | RUBIES-EGS-41280 | EGS63-30081 | RUBIES-UDS-170428 | RUBIES-UDS-170824 | RUBIES-UDS-140125 |
| ASTRODEEP ID | CEERS-43510 | CEERS-39492 | CEERS-47111 | CEERS-33796 | PRIMER-UDS-133345 | PRIMER-UDS-127017 | PRIMER-UDS-95456 |
| Other ID | o006_s00089aaAs reported in 2025ApJ...980..230T. | o006_s35616aaAs reported in 2025ApJ...980..230T. | o005_s41280aaAs reported in 2025ApJ...980..230T. | ||||
| CEERS-EGS-BD-4bbAs reported in 2024ApJ...964...66H. | |||||||
| R.A. (J2000) | 2149102828 | 214938816 | 214828434 | 214929872 | 3421396 | 34258098 | 34315054 |
| Decl. (J2000) | +52860077 | +52873855 | +52810819 | +52856079 | 510655 | 5106006 | 5149353 |
| NIRCam Photometry | |||||||
| F115W fν (nJy) | 71.12 | 9334 | 2243 | 212 | 226 | 247 | 1607 |
| F150W fν (nJy) | 383 | 13086 | 2263 | 214 | 285 | 86 | 1276 |
| F200W fν (nJy) | 272 | 15206 | 2333 | 233 | 304 | 65 | 1095 |
| F277W fν (nJy) | 13.31.9 | 10805 | 1542 | 163 | 243 | 24 | 693 |
| F356W fν (nJy) | 40.01.5 | 13863 | 2072 | 402 | 633 | 223 | 1534 |
| F444W fν (nJy) | 1143 | 11345 | 1823 | 383 | 794 | 835 | 3715 |
| mF115W (AB mag) | 26.770.06 | 23.970.05 | 25.530.05 | 28.110.13 | 28.060.28 | 27.90.3 | 25.890.07 |
| mF150W (AB mag) | 27.450.10 | 23.610.05 | 25.520.05 | 28.100.21 | 27.780.19 | 28.3 | 26.140.07 |
| mF200W (AB mag) | 27.810.11 | 23.450.05 | 25.480.05 | 27.990.16 | 27.710.15 | 28.5 | 26.310.07 |
| mF277W (AB mag) | 28.590.16 | 23.820.05 | 25.930.05 | 28.390.21 | 27.940.15 | 28.8 | 26.800.07 |
| mF356W (AB mag) | 27.390.06 | 23.550.05 | 25.610.05 | 27.390.07 | 26.900.07 | 28.040.18 | 25.940.06 |
| mF444W (AB mag) | 26.260.06 | 23.760.05 | 25.750.05 | 27.440.10 | 26.660.07 | 26.610.08 | 24.980.05 |
| NIRSpec/Prism Spectroscopy | |||||||
| 1.2 m S/N | 5 | 54 | 14 | 3.2 | 6 | 1.9 | 17 |
| 4.0 m S/N | 8 | 22 | 7 | 1.2 | 8 | 5 | 12 |
Our seven sources were selected among point sources identified in the CEERS/EGS and PRIMER/UDS NIRCam imaging , with spectral data grades 2.5 (a quality assessment metric indicating the reliability of spectral features) and automated redshift estimates of . These criteria initially narrowed our selection to six sources in the EGS field and three sources in the UDS field. We visually inspected NIRCam images to ensure the sources were consistent with point sources, and visually confirmed that the NIRSpec/Prism spectra were consistent with low-temperature dwarf morphologies as opposed to active galactic nuclei or other compact extragalactic sources. This evaluation allowed us to reject two duplicate EGS sources111, resulting in a final sample of seven low-temperature stars and brown dwarfs, summarized in Table 1. We list in the NIRCam F115W, F150W, F200W, F277W, F356W, and F444W filters from the CEERS/EGS and PRIMER/UDS surveys as compiled in the ASTRODEEP-JWST catalog (2024A&A...691A.240M). For the corresponding magnitudes, we conservatively include a systematic uncertainty of 0.05 mag to account for zeropoint errors. Three of the sources—RUBIES-BD-1, -2, and -3—were by 2024ApJ...964...66H and 2025ApJ...980..230T, whose results we discuss below in comparison with our findings.
RUBIES-BD-1 RUBIES-BD-2
RUBIES-BD-3 RUBIES-BD-4
RUBIES-BD-5 RUBIES-BD-6
RUBIES-BD-7
Figure 1 displays the NIRSpec/Prism spectra for these sources, flux-calibrated to their apparent F444W magnitudes and presented in flux density units.222 RUBIES-BD-1, -6, and -7 exhibit sharp flux peaks at 1.0, 1.25, 1.6, and 2.1 m, and a broad peak at 4 m, shaped by strong H2O and CH4 bands that are typical of T dwarfs (Burgasser_2006; 2012ApJ...760..151S; 2024ApJ...973..107B). RUBIES-BD-2 displays broader flux peaks at 1.2, 1.65, and 2.15 m between strong H2O absorption bands; and a blue, relatively featureless spectral slope toward longer wavelengths, typical of L dwarfs (2005ApJ...623.1115C; 2012ApJ...760..151S). RUBIES-BD-3, -4, and -5 have features intermediate between these morphological types (only marginally discernible in the spectrum of our faintest target RUBIES-BD-4), suggesting that they are late-L or early-T dwarfs. RUBIES-BD-3 also shows a peculiar split in the 1.2 m band and a bluer overall spectral morphology which may be associated with subsolar metallicity, as discussed below.
2.3 Classification and Distance Estimation
Spectral classifications for these sources were determined by comparing their 0.9–2.4 m spectra to L and T dwarf spectral standards defined in 2010ApJS..190..100K and Burgasser_2006, respectively. Near-infrared spectral data for these standards were drawn from the SpeX Prism Library Analysis Toolkit (2017ASInC..14....7B); see associated references in Table 7 of the Appendix. We also selected low-metallicity L and T dwarf spectral standards encompassing mild subdwarf (d/sd), subdwarf (sd), and extreme subdwarf (esd) classes as defined in 2017MNRAS.464.3040Z; 2018MNRAS.479.1383Z; 2019AJ....158..182G; and 2025ApJ...982...79B, drawing on published near-infrared spectral data acquired with various instruments for comparison . All spectra were smoothed and interpolated onto the same wavelength grid corresponding to the NIRSpec/Prism data. For our source spectra, we also applied small fractional pixel shifts of up to 1.5 pixels before interpolation, as determined by visual comparison to the best-fit standards, to account for variations in the projection of spectral data on the NIRSpec detector from the MSA slits.
For RUBIES-BD-1, -2, -3, -6, and -7, the best-fitting standard was identified by minimizing the statistic:
where is the observed flux density of the source, are the associated uncertainties, is the standard spectrum flux density, and is an optimal scaling factor that minimizes :
(cf. 2005ApJ...623.1115C). For RUBIES-BD-4 and -5, the spectral S/N is sufficiently low that the statistic cannot differentiate between adjacent standards, so we conducted a “by-eye” comparison to the spectral standards to identify an approximate best fit while allowing for generous uncertainties.
| Name | Unit | BD-1 | BD-2 | BD-3 | BD-4 | BD-5 | BD-6 | BD-7 |
| Classifications | ||||||||
| aaSignificant variance among the individual filter distance estimates. | (m) | 0.010 | 0.020 | 0.032 | +0.046 | 0.018 | 0 | 0.029 |
| SpT | T70.5 | L10.5 | sd:L61 | T23 | T12 | T81 | T50.5 | |
| Spectrophotometric Distance Estimates | ||||||||
| (pc) | 1000170 | 1320100 | 1870180aaSignificant variance among the individual filter distance estimates. | 2990230 | 2500250 | 940140 | 83090 | |
Note. — corresponds to the wavelength shift applied to the spectrum to align features with the best-fit standard. This shift is applied as a fractional pixel shift before spectra are interpolated onto a common wavelength grid, and accounts for variations in spectral projection on the NIRSpec detector by the MSA pixel mask.
Sources and their associated best-fit standards are shown in Figure 2 and summarized in Table 2. RUBIES-BD-1, -2, -6, and -7 are well-matched to T7, L1, T8, and T5 dwarf standards, respectively, and we estimate uncertainties of 0.5 subtypes for these sources. For RUBIES-BD-4 and -5, our best “by-eye” fits are to the T2 and T1 standards, respectively, with deviations generally consistent with the lower S/N of the spectra. For these sources, we estimate larger classification uncertainties of 3 and 2 subtypes, respectively. For RUBIES-BD-3, the minimum among the dwarf standards is achieved for the L3 dwarf, but there are significant in the 1.0–1.4 m -band region. We therefore compared this spectrum to the broader subdwarf grid (Figure 3), finding the d/sdL7 mild subdwarf and sdL5 subdwarf standards to be the best fits (minimum ). We adopt a classification of sd:L61 for this source.
With these classifications, we estimated distances to each source NIRCam photometry, using the spectral type/absolute magnitude conversions from 2024ApJ...962..177B based on the Sonora Bobcat models (2021ApJ...920...85M). For RUBIES-BD-1, -2, -4, -5, -6, and -7 we assumed a metallicity [M/H] = 0.0; for RUBIES-BD-3 we assumed [M/H] = 0.5. Uncertainties include contributions from the measured photometry and spectral classifications, and were combined using Monte Carlo techniques. Distance estimates are listed in Table 2, and are in agreement across all six filters for all of the sources with the exception of RUBIES-BD-3, which varies from pc for F115W photometry to pc for F356W photometry. This 3.6 discrepancy is likely associated with the metal-poor nature of the source. We computed uncertainty-weighted average distance estimates by using the higher of the lower and upper error bounds as the weighting factor (), and computed the combined uncertainty as
where = 3–6 is the number of filters with distance estimates. This expression accounts for both the scatter among distance estimates and individual distance uncertainties. Distances range from 83090 pc for RUBIES-BD-7 to nearly 3 kpc for RUBIES-BD-4, with five of the sources having estimated distances exceeding 1 kpc. At the Galactic latitude of the EGS and UDS fields, these distances correspond to vertical offsets from the disk , extending into the volume dominated by thick disk objects (2008ApJ...673..864J; 2024ApJ...962..177B).
3 Spectral Model Fits
3.1 Model Grids
To characterize the physical properties of these sources, we compared the NIRSpec spectra to three different model sets, allowing us to assess the quality of the fits across different model assumptions. Table 3 summarizes the models and associated parameter ranges explored. The Sonora Elf Owl models (2024ApJ...963...73M) are cloud-free radiative-convective equilibrium models that include the effects of disequilibrium chemistry due to vertical mixing, and allow for variations in temperature (Teff), surface gravity (), bulk metallicity ([M/H]), the C/O elemental abundance ratio, and the vertical diffusion coefficient (). We use the version of these models that have artificially suppressed PH3 abundances (see 2024ApJ...973...60B ). The Sonora Diamondback models (Morley_2024) are based on the same framework but include refractory clouds in both the chemistry and atmospheric opacity using the approach of 2001ApJ...556..872A. These models are computed in chemical equilibrium and include an parameter to account for the efficiency of condensate sedimentation. Here, we adopt a value of = 10 to denote models without clouds. The SAND (Spectral ANalog of Dwarfs) model set (2024ApJ...971...65G; 2024RNAAS...8..134A) is based on the PHOENIX code (1997ApJ...483..390H), and was designed to explore the non-solar abundances of brown dwarfs in globular clusters and the Milky Way’s thick disk and halo. This grid samples variations in alpha element enrichment ([/Fe]), accounts for self-consistent condensation and gravitational settling of clouds following 2012RSPTA.370.2765A, and assumes chemical equilibrium.
| Parameter | Unit | Elf Owl [1] | Diamondback [2] | SAND [3,4] | aaWidth of normal distribution for parameter updated in MCMC fitting. |
| Teff | (K) | 500 to 2400 | 900 to 2400 | 700 to 3000 | 35 |
| (cm s-2) | 4.5 to 5.5 | 4.5 to 5.5 | 4.5 to 6.0 | 0.2 | |
| (dex) | 0.5 to +0.5 | 0.5 to +0.5 | 0.35 to +0.3 | 0.2 | |
| (cm2 s-1) | 2.0 to 9.0 | 0.25 | |||
| C/ObbRelative the Sun, for which we adopt a value of C/O = 0.458 (2003ApJ...591.1220L). | 0.5 to 1.5 | 0.05 | |||
| ccA value of = 10 is adopted for Diamondback models without clouds; see Morley_2024. | 1 to 10 | 0.25 | |||
| (dex) | 0.05 to +0.40 | 0.1 |
References. — [1] 2024ApJ...963...73M; [2] Morley_2024; [3] 2024RNAAS...8..134A; [4] 2024ApJ...971...65G.
All data and models were interpolated onto a common wavelength scale that matches the variable resolution and wavelength sampling of the NIRSpec/Prism mode, and were evaluated in flux density units scaled to each source’s apparent F444W magnitudes. Because the models are computed in surface flux densities, the optimal scaling factory between data and models (Eqn. 2) is equal to the ratio , where is the radius of the source and its distance. We assumed a fixed radius of 0.8 RJup, expected for an evolved, relatively massive brown dwarf.333 (2021ApJ...920...85M) scale factor provides a second estimate of the source distance = 1.8510 pc.
3.2 Markov Chain Monte Carlo Fitting
We used the ucdmcmc package (ucdmcmc) to fit the three model sets to our brown dwarf spectra. Spectra and models were compared using the same chi-square statistic as Eqn. 1. A first fit was conducted to identify the best individual model within a model set based on the minimum value. These model parameters were then used to initiate a Metropolis-Hastings Markov Chain Monte Carlo (MCMC) algorithm (1953JChPh..21.1087M; HASTINGS01041970) to determine the overall best-fit parameters and their uncertainties from the posterior distributions. Model logarithmic flux densities were linearly interpolated between grid points over logarithmic parameter values (e.g., interpolates as log ). The MCMC algorithm proposes a new set of parameters by varying the previous set using normal distributions with the fixed widths listed in Table 3. The difference between the values of the proposed () and previous () parameters were compared to the overall minimum , which in turn was compared to a random number drawn from a uniform distribution between 0 and a threshold value :
If this condition was satisfied, the proposed parameters were adopted; otherwise, the current parameters were retained. Through experimentation, we found an optimal value of = 0.1. The MCMC algorithm proceeds over 5,000 steps in a single chain, and the first 25% of the chain was removed for burn-in. Convergence was assessed based on the consistency of the chains and the existence of single optimal values for well-fit parameters in the posterior distributions. Parameters and their uncertainties were computed as weighted quantiles of the marginalized distribution across the chain, using the weighting function:
which accounts for the degrees of freedom = (740 spectral data points) - (5–6 model parameters) for each model fit. We adopted the 50% quantile (median) as the fit value and the 16% and 84% quantiles as estimates of the lower and upper uncertainty bounds. For each set of models, we conducted separate fits for both the 0.9–2.4 m near-infrared (NIR) range and the full 0.9–5.1 m range of the NIRSpec/Prism data.
4 Results
Best fit models for each source and model set are displayed in Figures 4 and 5 and fit parameters are listed in Tables 4 and 5 for the 0.9–2.4 m (NIR) and 0.9–5.1 m (NIRSpec) regions, respectively. A summary of fit parameters across models is provided in Table 6, while a full list of parameter values and distribution visualizations are provided in a separate Zenodo link.444 Here, we describe the outcomes of the fits and the physical interpretation of the fit parameters for individual sources.
4.1 RUBIES-BD-1
The best NIR and NIRSpec fits for this T7 dwarf are from the Elf Owl and Diamondback models, with a slight preference for the former with modest vertical mixing ( = 3.11.0 averaged between NIR and NIRSpec fits). The Diamondback models require very efficient sedimentation or no clouds ( = 9.50.4). Averaging the results of the NIRSpec fits for these models yields Teff = 930 K, = 4.90.4, and [M/H] = +0.1, Elf Owl models a solar C/O ratio. The inferred effective temperature is slightly higher but consistent with previously established trends in spectral type (e.g., Teff = 830130 K for a T70.5 dwarf based on 2015ApJ...810..158F). This modest overshoot may be due to the best-fitting models failing to match the depth of the 1.6 m CH4 band or the strong CO2 band at 4.1 m (see discussion below). Nevertheless, the agreement in classification and model fit parameters indicate that this is a normal late T-type brown dwarf.
RUBIES-BD-1 was as CEERS-EGS-BD-4 in 2024ApJ...964...66H, identified from multi-band JWST/NIRCam photometry. The photometric measurements were fit to three sets of atmosphere models: Sonora Bobcat and Sonora Cholla (2021ApJ...923..269K), precursors to Diamondback and Elf Owl models; and ATMO (2020A&A...637A..38P). The fits yield Teff = 1000–1050 K and = 5.0, but no evaluation of metallicity. The authors infer a distance of 1350–1831 pc for estimated radii of 0.92–1.2 RJup, somewhat further than our estimate of 1030140 pc using a smaller radius estimate. RUBIES JWST/NIRSpec spectral data of this source were reported in 2025ApJ...980..230T, where the source is identified as o006_s00089. These authors compared the NIRSpec spectrum to Elf Owl and ATMO 2020++ (2023AJ....166...57M) model sets, the latter an updated version of the ATMO models. For the , the authors inferred a similar Teff = 955 K but a considerably smaller = 4.34, outside the range considered in these fits, albeit formally consistent with our uncertainties. also infer a much larger distance for this source, 2.00.2 kpc, by assuming a radius of 1.2 RJup as suggested by the low surface gravity, but estimate a shorter distance of 1.4–1.7 kpc if the radius is constrained to 0.8–0.9 RJup. The authors find a significantly supersolar metallicity of [M/H] = +0.410.05 from their model fits, above but again formally consistent with our fits. 2025ApJ...980..230T note that their parameter uncertainties, which are considerably smaller than those reported here, may be underestimated.
4.2 RUBIES-BD-2
The NIR spectrum for this L1 dwarf is best fit by the Diamondback models, while the full NIRSpec spectrum has equal values between the Diamondback and SAND models. As both models include a prescription for cloud condensation and scattering, it is likely the presence of clouds that rules out the condensate-free Elf Owl models as a reliable fit. Indeed, Diamondback model fits to both the NIR and NIRSpec spectra require relatively thick clouds ( = 1.81.0). However, despite providing the best fits, Figure 5 shows that the Diamondback and SAND models have difficulty reproducing the 0.9–2.4 m spectral slope of RUBIES-BD-2, the former underestimating the 1.6 and 2.1 m flux peaks and the latter overestimating the 1.2 and 1.6 m peaks. We also find that the fits between the NIR and NIRSpec ranges yield marginally distinct temperatures, with Teff = 1760 K for the former (Diamondback only) and Teff = 2040 K for the latter (average of Diamondback and SAND), a 1.7 difference. The higher temperature is more in line with local L10.5 dwarfs (Teff = 2100130 K from 2015ApJ...810..158F). We also find a 3 difference in the inferred value between the SAND models ( = 4.22) and Diamondback models ( = 5.30) based on fits to the NIRSpec data; and a 4 discrepancy between our model-scaled distance estimate of 860 pc and our spectrophotometric distance estimate of 1320100 pc. These deviations suggest that the Elf Owl, Diamondback, and SAND models all have difficulty reproducing the spectrum of RUBIES-BD-2, indicating potentially missing elements to the models.
RUBIES JWST/NIRSpec spectral data of this source were in 2025ApJ...980..230T, where the source is identified as o006_s35616. The authors compared these data to Elf Owl and BT-Settl CIFIST models (2012RSPTA.370.2765A; 2011SoPh..268..255C), finding Teff = 1810–1910 K, lower but consistent with our NIRSpec fits. However, the authors again infer a very low = 3.250.01 and a supersolar [M/H] = +1.000.01 from their Elf Owl fits, inconsistent with our findings. 2025ApJ...980..230T note significant residuals in their model fits, recognizing that neither the Elf Owl nor BT-Settl models include condensate opacity. They introduce a reddening parameter based on the interstellar extinction law of 2024ApJ...964L...3W, and find that = 3 mag substantially improves their Elf Owl model fits while yielding a substantially higher Teff = 2286 K and subsolar [M/H] = 0.710.02. It is likely that a combination of atmospheric condensate opacity and interstellar reddening are required to accurately reproduce these data.
4.3 RUBIES-BD-3
As discussed in Section 2, RUBIES-BD-3 is better matched to subdwarf standards, suggesting that it should be well-matched to metal-poor atmosphere models. However, both NIR and NIRSpec fits yield consistently better fits to supersolar metallicity models. In the NIR range, Elf Owl, Diamondback, and SAND models all yield comparable best-fit values; in the NIRSpec range, the Elf Owl models are somewhat less reliable. Primary fit parameters of Teff, , and [M/H] are in agreement across all three models and fit ranges, with a relatively high Teff = 2170130 K from the NIRSpec fits, notably warmer than the earlier-type L dwarf RUBIES-BD-2. Secondary parameters are similar to NIRSpec fits for RUBIES-BD-2, with all three models indicating relatively low surface gravities ( = 4.70.4), and for the Diamondback models thick clouds ( = 2.1). The Elf Owl model fits also indicate a high C/O = 0.73 and modest vertical mixing ( = 3.6). However, like RUBIES-BD-2, all three models have clear flaws in reproducing the observed spectrum, particularly in the NIR range where Elf Owl models show excess flux shortward of 1.3 m, and SAND models fail to match the shape of the 1.6 m peak. None of the models accurately reproduce the 1.2 m “notch” absorption feature that is characteristic of L- and T-type subdwarfs (Figure 3). Again, the models deployed for this seemingly metal-poor L dwarf may be missing a critical element of atmospheric chemistry, composition, or dynamics that may be skewing the fits toward higher metallicities.
2025ApJ...980..230T reported RUBIES JWST/NIRSpec spectral data for this source, identified as o005_s41280, and compared the data to Elf Owl and BT-Settl models. The authors find a similar Teff = 2160–2260 K from these fits without taking into account reddening, with an approximately 200 K increase when a modest reddening of = 1.0-1.5 mag is taken into account. In both cases, the authors find a very low surface gravity for this source, = 3.3–3.5, well outside our fitting parameter range. As with RUBIES-BD-2, such a low surface gravity is highly unlikely for a field brown dwarf as it would require an age 1 Myr (2023A&A...671A.119C).
4.4 RUBIES-BD-4
The relatively low S/N of the RUBIES-BD-4 spectrum results in equivalent matches among all three models in both NIR and NIRSpec ranges, and relatively large uncertainties in inferred parameters. The poorly-constrained Teff = 1800 K is higher but not inconsistent with trends among solar metallicity dwarfs for its T23 classification (Teff = 1180200 K; 2015ApJ...810..158F). Notably, SAND model fits to the NIRSpec data yield a considerably lower Teff = 1050 K as compared to the other models, even though the best-fit model has a Teff = 1506 K (Figure 5). Secondary parameters indicate efficient vertical mixing ( = 7.81.0 from Elf Owl models), thin clouds ( = 8.40.7 from Diamondback models), and near-solar elemental abundances ([M/H] = 0.1 from all models, C/O = 0.54 from Elf Owl models, and [/M] = +0.17 from SAND models). The large range in temperatures across the models results in a comparably large uncertainty in the estimated distance, from 2000 pc for SAND to 6400 pc for Diamondback, which straddle the uncertain 5400 pc distance estimate inferred from its spectral type and NIRCam photometry (Table 2).
4.5 RUBIES-BD-5
RUBIES-BD-5 is best fit in both the NIR and NIRspec ranges by the SAND models, yielding an atmospheric temperature Teff = 1270 K that is in line with its T12 classification (Teff = 1220150 K; 2015ApJ...810..158F). Figure 5 shows that the SAND models do a better job reproducing the 1–2.2 m and 4.0 m peaks compared to the Elf Owl and Diamondback models, both of which produce overly blue spectra. The NIRSpec SAND fits also yield a significantly subsolar metallicity ([M/H] = 0.7) and modest alpha enrichment ([/M] = +0.19), both indicative of a thick disk or halo origin.
4.6 RUBIES-BD-6
RUBIES-BD-6, like RUBIES-BD-4, has a relatively low S/N spectrum and is equally-well fit by all three models in both NIR and NIRSpec ranges, with a slight preference for the Elf Owl models in the latter. The NIRSpec fits yield an average Teff = 830130 K, high but consistent with its T81 classification (Teff = 680190 K; 2015ApJ...810..158F). Given the large uncertainties, there is no evidence of a metallicity or elemental compositions significantly distinct from solar, while the Diamondback models require efficient sedimentation ( = 8.0) as is typical for late T dwarfs.
4.7 RUBIES-BD-7
RUBIES-BD-7 is best fit in both the NIR and NIRSpec ranges by the Elf Owl and Diamondback models, the latter requiring very efficient sedimentation ( = 9.60.4, consistent with no clouds). The inferred Teff = 118080 K is consistent with other T50.5 dwarfs (Teff = 1030120 K; 2015ApJ...810..158F), while the metallicity ([M/H] = 0.20.3) and Elf Owl C/O = 0.48 are consistent with solar composition. The excellent agreement between the model scaling and spectrophotometric distances further indicate that this is likely a normal thin disk brown dwarf.
| Model | Teff | [] | C/O | [] | Min. | ||||
| (K) | (cm s-2) | (dex) | (cm2 s-1) | (dex) | (pc) | ||||
| RUBIES-BD-1 (T70.5) | |||||||||
| E | 2.8 | ||||||||
| D | 3.0 | ||||||||
| S | 3.5 | ||||||||
| RUBIES-BD-2 (L10.5) | |||||||||
| E | 77 | ||||||||
| D | 19 | ||||||||
| S | 183030 | 65020 | 29 | ||||||
| RUBIES-BD-3 (sd:L61) | |||||||||
| E | 3.8 | ||||||||
| D | 4.0 | ||||||||
| S | 4.2 | ||||||||
| RUBIES-BD-4 (T23) | |||||||||
| E | 1870360 | 1.8 | |||||||
| D | 1.7 | ||||||||
| S | 1.7 | ||||||||
| RUBIES-BD-5 (T12) | |||||||||
| E | 2.1 | ||||||||
| D | 2.1 | ||||||||
| S | 1.6 | ||||||||
| RUBIES-BD-6 (T80.5) | |||||||||
| E | 1.0 | ||||||||
| D | 1.1 | ||||||||
| S | 1.0 | ||||||||
| RUBIES-BD-7 (T50.5) | |||||||||
| E | 114080 | 710100 | 13 | ||||||
| D | 1219100 | 850150 | 15 | ||||||
| S | 18 | ||||||||
Note. — Model labels are: E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A).
| Model | Teff | [] | C/O | [] | Min. | ||||
| (K) | (cm s-2) | (dex) | (cm2 s-1) | (dex) | (pc) | ||||
| RUBIES-BD-1 (T70.5) | |||||||||
| E | 89080 | 980170 | 2.6 | ||||||
| D | 3.1 | ||||||||
| S | 4.0 | ||||||||
| RUBIES-BD-2 (L10.5) | |||||||||
| E | 49 | ||||||||
| D | 24 | ||||||||
| S | 205050 | 0.05aaMCMC fits were unable to converge on this parameter. | 24 | ||||||
| RUBIES-BD-3 (sd:L61) | |||||||||
| E | 3.3 | ||||||||
| D | 2.8 | ||||||||
| S | 2.7 | ||||||||
| RUBIES-BD-4 (T23) | |||||||||
| E | 2.2 | ||||||||
| D | 2.3 | ||||||||
| S | 2.1 | ||||||||
| RUBIES-BD-5 (T12) | |||||||||
| E | 3.7 | ||||||||
| D | 3.0 | ||||||||
| S | 1.6 | ||||||||
| RUBIES-BD-6 (T80.5) | |||||||||
| E | 1.1 | ||||||||
| D | 1.4 | ||||||||
| S | 1.3 | ||||||||
| RUBIES-BD-7 (T50.5) | |||||||||
| E | 117080 | 7 | |||||||
| D | 118690 | 8 | |||||||
| S | 0.05aaMCMC fits were unable to converge on this parameter. | 14 | |||||||
Note. — Model labels are: E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A).
| Parameter | Unit | BD-1 | BD-2 | BD-3 | BD-4 | BD-5 | BD-6 | BD-7 |
| SpT | T70.5 | L10.5 | sd:L61 | T23 | T12 | T80.5 | T50.5 | |
| Best Model(s) | E(D) | DS | (E)DS | EDS | S | E(DS) | E(D) | |
| Teff | (K) | 2170130 | 830130 | 118080 | ||||
| (cm s-2) | 4.90.4 | aaThe SAND NIRSpec fit included in this average has a significantly discrepant value of . | 5.20.5 | 4.90.4 | ||||
| [M/H] | (dex) | +0.10.3 | +0.20.4 | 0.20.3 | ||||
| (pc) | 2500300 | |||||||
| (pc) | 1000170 | 1320100 | 1870180 | 2990230 | 2500250 | 940140 | 83090 | |
| (pc) | 1030140 | 980210bbModel and spectrophotometric distance estimates differ significantly for this source. | 2000300 | 3000400 | 2300300 | 980170 | 81070 | |
| (pc) | 890120 | 850180 | 1700300 | 2990200 | 2200200 | 810210 | 72080 | |
| (thin disk) | 507% | 5210% | 137% | 2.32.2% | 95% | 568% | 643% | |
| (thick disk) | 455% | 438% | 683% | 626% | 67.91.7% | 407% | 333% | |
| (halo disk) | 5.11.2% | 4.81.2% | 185% | 357% | 235% | 4.41.4% | 3.00.4% |
Note. — Best model(s) are for NIRSpec fits, and are indicated by labels E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A). When more than one model is listed, models outside of parentheses have indistinguishable best-fit , while models in parentheses have slightly higher (more poorly fit) best-fit . Model parameters are determined from an uncertainty-weighted mean of the posterior distributions of all the best-fit models listed (see Eqns.5 and 3). The adopted distance is an uncertainty-weighted average of the model fit distance () and spectrophotometric distance () assuming the larger of the error bounds.
5 Discussion
5.1 Population Assignments
The large distances and high Galactic latitudes of these sources translate into large vertical scaleheights from the plane, giving them a high probability of belonging to the Milky Way’s thick disk or halo populations (2024ApJ...962..177B). To quantify population membership, we used the star density model of 2008ApJ...673..864J to estimate the relative space densities of thin disk, thick disk, and halo sources in the directions and at the distances for each of our targets, sampling distance uncertainties using Monte Carlo methods. The resulting probabilities are listed in Table 6. Not surprisingly, the most distant sources in our sample, RUBIES-BD-3, -4, and -5, have high probabilities of thick disk membership, with RUBIES-BD-4 a significant probability (357%) of halo membership. This subsample includes both the classified L subdwarf RUBIES-BD-3 and the only source in the sample whose modeled parameters indicate subsolar metallicity, RUBIES-BD-5, The mid- to late-T dwarfs in our sample—RUBIES-BD-1, -6, and -7—are intrinsically fainter and detected at vertical scaleheights 1 kpc, and are thus more likely to be thin disk dwarfs, as supported by their spectral similarity to local spectral standards. The mix of thin disk, thick disk, and halo brown dwarfs mimics the composition of distant stars and brown dwarfs detected in other deep JWST survey samples (2024ApJ...962..177B; 2024ApJ...964...66H; 2024ApJ...975...31H; 2025arXiv251000111H).
5.2 Spectral Variations Among Distant T Dwarfs
The mid- and late-type T dwarfs that have been identified in deep JWST/NIRSpec spectral surveys extend out to 1–2 kpc, a distance that encompasses a diversity of populations, ages, and compositions. As such, these early discoveries provide the opportunity to examine the influence of secondary parameters such as surface gravity and metallicity on cool brown dwarf spectra. To illustrate this, Figure 7 compares the NIRSpec/Prism spectra of five distant mid- and late-type T dwarfs, including RUBIES-BD-1 (T7), -6 (T8), and -7 (T5), the T6 dwarf UNCOVER-BD-2 (estimated distance of 2.1–2.6 kpc; 2024ApJ...962..177B), and the T6 dwarf JADES-GS-BD-9 (estimated distance of 1.8–2.3 kpc; 2024ApJ...975...31H). All of these sources are similarly normalized to focus on spectral morphology, and are compared to the NIRSpec/Prism spectrum of the T6 dwarf standard SDSS J1624+0029 (2024ApJ...973..107B). We separately evaluate the 0.9–2.4 m NIR and 3.6–5.0 m MIR regions to focus on key spectral features. In the NIR, we observe subtle variations between these spectra in several key features, including the depth of the 1.15 m and 1.6 m CH4 bands (shallowest in RUBIES-BD-7, deepest in RUBIES-BD-1), the width of the 1.25 m peak (widest in RUBIES-BD-7, narrowest in JADES-GS-BD-9 and RUBIES-BD-6), and the shape and brightness of the 2.1 m peak (broad in RUBIES-BD-7, shifted redward in UNCOVER-BD-2, shifted blueward in RUBIES-BD-1, flat in JADES-GS-BD-9 and RUBIES-BD-6). In the MIR, the broad 4 m peak shows a variety of shapes among these sources, from a relatively featureless and blueshifted peak in the spectra of RUBIES-BD-7 and UNCOVER-BD-2, to the more structured peaks in the spectra of JADES-GS-BD-9 and RUBIES-BD-7, the latter showing strong CO2 and CO absorption bands at 4.3 m and 4.6 m, also observed in the spectrum of SDSS J1624+0029.
To contextualize these variations, we compare in the right panels of Figure 7 Elf Owl atmosphere models with variations in temperature, surface gravity, metallicity, and vertical mixing efficiency relative to a baseline model appropriate for a thin disk T6 dwarf (Teff = 800 K, = 4.5, [M/H] = 0, and = 4). NIR variations sensitive to these features have been discussed extensively in the literature (e.g., 2002ApJ...564..421B; 2010ApJ...710.1627L; 2012MNRAS.422.1922P; 2017MNRAS.468..261Z; 2024ApJ...963...73M; 2025ApJ...982...79B). Here, we highlight that variations in the depths of CH4 bands, the width of the 1.25 m peak, and shape of the 2.1 m peak are all driven primarily by temperature and metallicity variations in the models. Surface gravity plays a greater role in setting the relative brightness of the 1.05 m and 1.6 m peaks, while vertical mixing plays only a minor role in shaping the 2.1 m peak. In the MIR, the models show that metallicity plays greatest role in shaping the 4 m peak, with low metallicity models missing the CO2 and CO bands and producing a relatively featureless broad plateau. Vertical mixing also modulates the CO2 feature. Temperature primarily shifts the location of the 4 m peak (blueward for warmer temperatures), while surface gravity only modestly influences the depth of the CO band.
This analysis suggests that all of these parameters can drive the variations observed among these similarly-classified distant T dwarfs. As such, a comparison of multiple features is necessary to isolate metallicity variations from other secondary parameters. There is an important caveat to this analysis, in that current models do not fully capture the complex atmosphere chemistry among the observed molecular constituents, Most importantly, the 4.3 m CO2 band is known to be far stronger than can be explained even with efficient vertical mixing (2010ApJ...722..682Y; 2024ApJ...973...60B). We also note that none of these spectra show clear evidence of PH3 absorption that was initially claimed in the distant (0.6–1.2 kpc) late T dwarf UNCOVER-BD-3 (2024ApJ...962..177B, see 2024ApJ...973...60B for further discussion) and more recently detected in the metal-poor T dwarf companion Wolf 1130C (burgasser-ph3). The absence or weakness of this feature in most brown dwarf and exoplanet spectra a challenge to chemical models (2018ApJ...858...97M; 2024ApJ...977L..49R).
6 Summary
The key results of this study are as follows:
We have identified seven brown dwarfs observed with JWST/NIRSpec as part of the RUBIES survey. All seven sources show spectral features characteristic of L- and T-type brown dwarfs in the 1–5 m range, including strong H2O, CH4, CO, and CO2 molecular features.
We determined spectral types for these sources by comparing their near-infrared (0.9–2.4 m) spectra to dwarf and subdwarf standards, inferring classifications spanning L1 to T8. One source, RUBIES-BD-3, is best matched to mild L subdwarf (d/sdL) and L subdwarf (sdL) standards, suggesting that is has a metal-poor atmosphere.
We compared flux-calibrated spectra to three sets of atmosphere models: Sonora Elf Owl, Sonora Diamondback, and SAND. We find the L dwarf spectra are best matched to the Diamondback and SAND models, likely due to the presence of condensates in their atmospheres; the early-T dwarf spectra are best matched to the SAND models; and the mid- and late-T dwarf spectra are best matched to the Elf Owl and Diamondback models, the latter with efficient sedimentation or no clouds. Spectra of the two L dwarfs, RUBIES-BD-2 and -3, show the most significant deviations from the models, possibly due to incorrect treatment of condensates or the failure to incorporate reddening corrections, as previously noted by 2025ApJ...980..230T. Only one source, RUBIES-BD-5, shows significant evidence of being metal-poor based on the model fits.
Three of our sources—RUBIES-BD-3, -4, and -5—have estimated vertical scaleheights exceeding 1 kpc, and are mostly likely members of the Galactic thick disk population. RUBIES-BD-4, our faintest, most distant, and most poorly characterized source, also has a significantly probability of being a member of the Galactic halo population.
Examination of the present sample of mid- and late-type T dwarfs found in deep JWST spectral surveys at distances of 1–2 kpc show subtle but distinct variations in the strengths of major molecular bands and the shapes of inter-band peaks, most notably the 2.1 m flux peak and the 4.3 m CO2 band. Examination of comparable Elf Owl models point to multiple factors that may be responsible for these variations, requiring further examination to isolate the these distant objects.
This study adds to the growing sample of deep brown dwarf spectra from JWST/NIRSpec Prism data collected in deep survey programs such as CEERS (2023ApJ...946L..13F), UNCOVER (2024ApJ...974...92B), JADES (2025ApJS..277....4D), CANUCS (2022PASP..134b5002W), BoRG (2025ApJ...983...18R), NEXUS (2024arXiv240812713S), and CAPERS (2024jwst.prop.6368D), among others (cf. 2025arXiv251002026T). these discoveries increase the diversity of physical parameters sampled among known brown dwarfs, and provide the opportunity to improve atmosphere and evolutionary models of low-temperature atmospheres more broadly, not only for thick disk and halo brown dwarf but also candidate brown dwarfs identified in globular clusters (2024ApJ...971...65G) and stellar stream populations (2025ApJ...982...79B), as well as gas giant exoplanets orbiting metal-poor stars (2019A&A...621A.110B; 2022MNRAS.517.4447A). Building larger and more diverse samples of cool brown dwarfs from these surveys can also facilitate their use to address broader questions in Milky Way science, such as the evolution and distribution of chemical abundances, the Milky Way’s star formation and accretion history, and the dynamical history of various Galactic populations (2004ApJS..155..191B; 2017ApJ...847...53R; 2022ApJ...934...73A; 2024ApJ...967..115B; 2024ApJ...974..222R; 2025ApJ...985...48H).
Appendix A Near-Infrared Spectral Standards
Table 7 lists the near-infrared spectral standards used for the classification of our sources in Section 2.3. Dwarf standards are defined in Burgasser_2006 and 2010ApJS..190..100K; subdwarf standards are defined in 2017MNRAS.464.3040Z; 2018MNRAS.479.1383Z; 2019AJ....158..182G; and 2025ApJ...982...79B. We drew from previously-published data acquired with several different instruments, which were resampled to the wavelength-dependent resolution of the JWST/NIRSpec prism mode (Birkmann_2022). The resampled spectra are included as part of the online materials for the article.
| SpT | Source | Instrument | Data Ref. |
| L0 | 2MASP J0345432+254023 | IRTF/SpeX | [1] |
| L1 | 2MASSW J2130446084520 | IRTF/SpeX | [2] |
| L2 | Kelu1AB | IRTF/SpeX | [3] |
| L3 | 2MASSW J1506544+132106 | IRTF/SpeX | [4] |
| L4 | 2MASS J215804571550098 | IRTF/SpeX | [2] |
| L5 | SDSS J083506.16+195304.4 | IRTF/SpeX | [5] |
| L6 | 2MASSI J1010148040649 | IRTF/SpeX | [6] |
| L7 | 2MASSI J0103320+193536 | IRTF/SpeX | [2] |
| L8 | 2MASSW J1632291+190441 | IRTF/SpeX | [4] |
| L9 | DENIS-P J02554700 | IRTF/SpeX | [7] |
| T0 | SDSS J120747.17+024424.8 | IRTF/SpeX | [8] |
| T1 | SDSSp J083717.22000018.3 | IRTF/SpeX | [9] |
| T2 | SDSSp J125453.90012247.4 | IRTF/SpeX | [10] |
| T3 | 2MASS J120956131004008AB | IRTF/SpeX | [10] |
| T4 | 2MASSI J2254188+312349 | IRTF/SpeX | [10] |
| T5 | 2MASS J15031961+2525196 | IRTF/SpeX | [10] |
| T6 | SDSSp J162414.37+002915.6 | IRTF/SpeX | [11] |
| T7 | 2MASSI J0727182+171001 | IRTF/SpeX | [11] |
| T8 | 2MASSI J0415195093506 | IRTF/SpeX | [10] |
| T9 | UGPS J072227.51054031.2 | IRTF/SpeX | [9] |
| d/sdL0 | 2MASS J00412179+3547133 | IRTF/SpeX | [10] |
| d/sdL1 | 2MASS J17561080+2815238 | IRTF/SpeX?? | [12] |
| d/sdL6 | SDSS J133148.92011651.4 | IRTF/SpeX | [2] |
| d/sdL8 | 2MASS J11582077+0435014 | IRTF/SpeX | [12] |
| d/sdL9 | CWISE J202130.11+152418.3 | Keck/NIRES | [13] |
| d/sdT0 | 2MASS J064531536646120 | VLT/Xshooter | [14] |
| d/sdT1 | WISE J030119.47231921.8 | IRTF/SpeX | [15] |
| d/sdT2 | WISE J000458.59133655.2 | IRTF/SpeX | [15] |
| d/sdT2.5 | CWISE J211255.59+303037.6 | Keck/NIRES | [13] |
| d/sdT4 | ULAS J002135.97+155226.8 | VLT/Xshooter | [16] |
| d/sdT5.5 | CWISE J113010.07+313944.7 | Keck/NIRES | [13] |
| d/sdT5.5 | GJ 576B | Keck/NIRES | [13] |
| d/sdT6 | CWISE J201342.30032643.4 | Keck/NIRES | [13] |
| d/sdT7.5 | LHS 6176B | Keck/NIRES | [13] |
| d/sdT8 | 2MASS J093935482448279 | IRTF/SpeX | [7] |
| d/sdT9 | ULAS J083338.11+005206.2 | Gemini/GNIRS | [17] |
| sdL0 | WISE J04592121+1540592 | IRTF/SpeX | [12] |
| sdL1 | ULAS J124947.04+095019.8 | VLT/Xshooter | [18] |
| sdL4 | ULAS J021642.97+004005.6 | VLT/Xshooter | [18] |
| sdL5 | SDSS J141624.08+134826.7 | IRTF/SpeX | [19] |
| sdT0 | WISEA J152443.14262001.8 | Keck/NIRES | [13] |
| sdT1 | CWISE J211250.11052925.2 | Keck/NIRES | [13] |
| sdT3 | CWISE J062316.19+071505.6 | Keck/NIRES | [13] |
| sdT4 | CWISE J155349.96+693355.2 | Keck/NIRES | [20] |
| sdT5.5 | CWISE J113019.19115811.3 | Keck/NIRES | [21] |
| sdT6 | 2MASS J09373487+2931409 | IRTF/SpeX | [11] |
| sdT7 | ULAS J141623.94+134836.3 | IRTF/SpeX | [22] |
| sdT7.5 | ULAS J001354.85+063445.4 | Gemini/GNIRS | [17] |
| esdL0 | SSSPM J101307341356204 | IRTF/SpeX | [23] |
| esdL1 | SSSPM J144420.67201922.2 | IRTF/SpeX | [2] |
| esdL3 | SDSS J125637.16022452.2 | IRTF/SpeX | [24] |
| esdL4 | 2MASS J16262034+3925190 | IRTF/SpeX | [23] |
| esdL6 | 2MASS J061643.51640719.8 | VLT/Xshooter | [18] |
| esdL7 | 2MASS J053253.46+824646.5 | Keck/LRIS+NIRES | [13] |
| esdT3 | WISEA J181006.18101000.5 | GTC/OSIRIS+Palomar/TSpec | [25,26] |
| esdT6 | WISEA J041451.67585456.7 | Magellan/FIRE | [25] |
References. — [1] 2006AJ....131.1007B; [2] 2014ApJ...794..143B; [3] 2007ApJ...658..557B; [4] 2007ApJ...659..655B; [5] 2006AJ....131.2722C; [6] 2006AJ....132..891R; [7] 2006ApJ...637.1067B; [8] 2007AJ....134.1162L; [9] 2017ASInC..14....7B; [10] 2004AJ....127.2856B; [11] 2006ApJ...639.1095B; [12] 2010ApJS..190..100K; [13] 2025ApJ...982...79B; [14] 2018MNRAS.480.5447Z; [15] 2019AJ....158..182G; [16] 2019MNRAS.486.1260Z; [17] 2014MNRAS.437.1009P; [18] 2017MNRAS.464.3040Z; [19] 2010AJ....139.1045S; [20] 2021ApJ...915..120M; [21] 2021ApJS..253....7K; [22] 2010AJ....139.2448B; [23] 2004ApJ...614L..73B; [24] 2009ApJ...697..148B; [25] 2020ApJ...898...77S; [26] 2022AA...663A..84L.